# Constrained Denoising, Empirical Bayes, and Optimal Transport

### v1

@@ Appendix D Proofs from Section 3
$\begin{cases}\textnormal{minimize}&\textnormal{Var}(\delta(Z)-\delta_{\mathcal{B}}(Z))\\ \textnormal{over}&\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}\\ \textnormal{with}&\mathbb{E}[\delta(Z)]=\mathbb{E}[\Theta]\\ \textnormal{and}&\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta).\end{cases}$ (D.1)
Now we apply a “Gaussianization trick” whereby we show that problem (D.1) has the same optimal value as the following problem:
$\begin{cases}\textnormal{minimize}&\textnormal{Var}(D-E)\\ \textnormal{over}&\textnormal{jointly Gaussian }(D,E)\\ \textnormal{with}&\textnormal{marginal distribution }D\sim\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\Theta))\\ \textnormal{and}&\textnormal{marginal distribution }E\sim\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))\\ \end{cases}$ (D.2)
To see that (D.2) $\leq$ (D.1), it suffices to find, for each $\delta$ that is feasible for (D.1), some $(D,E)$ that is feasible for (D.2) such that $\textnormal{Var}(D-E)=\textnormal{Var}(\delta(Z)-\delta_{\mathcal{B}}(Z))$. Of course, we can let
$\begin{pmatrix}D\\ E\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mathbb{E}[\Theta]\\ \mathbb{E}[\Theta]\end{pmatrix},\begin{pmatrix}\textnormal{Cov}(\Theta)&\textnormal{Cov}(\delta(Z),\delta_{\mathcal{B}}(Z))\\ \textnormal{Cov}(\delta_{\mathcal{B}}(Z),\delta(Z))&\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\end{pmatrix}\right)$ (D.3)
@@
To see that (D.1) $\leq$ (D.2), it suffices to show that the optimal value of (D.2) is achieved by a coupling which corresponds to some feasible $\delta$ for (D.1) (via (D.3)). When combined with the above, this will show that such a $\delta$ is optimal for (D.1).
Note that (D.2) is indeed the (squared) 2-Wasserstein distance between the two multivariate normal distributions $\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))$ and $\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\Theta))$. Indeed, recall [Dowson and Landau, 1982, Olkin and Pukelsheim, 1982] that the optimal value of (D.2) is exactly given by the Bures-Wasserstein distance
@@
so, in particular we can use (PD) to define
@@
Proof of Theorem 3.12.
First let us show that the integrand of (3.8) is well-defined for every $\pi\in\Gamma(F;G,\Psi)$. To see this, write $H$ for the second marginal of $\pi$, note that
@@
Next let us show that (3.8) admits solutions. By (C), there exists a Borel set $A\subseteq\mathbb{R}^{d}$ with $F(A)=1$ such that $c_{G}:A\times\mathbb{R}^{m}\to[0,\infty)$ is continuous. Since
@@
for all $\pi\in\Gamma(F;G)$, it follows from the Portmanteau lemma that the objective of (3.8) is weakly lower semi-continuous. Since a lower semi-continuous function on a compact set must achieve a minimizer, it only remains to show that the feasible set $\Gamma(F;G,\Psi)$ is weakly compact. To do this, note that we can write $\Gamma(F;G,\Psi)=\bigcup_{H\in\mathcal{H}}\Gamma(F;H)$ for
@@
as $r\to\infty$. This proves that $\mathcal{H}$ is tight, and hence guarantees the existence of a solution to (3.8).
Now we finish the proof, by applying the result in García-Trillos and Sen [2024, Theorem 2.4 and Remark 4.2] which shows that, under assumptions (BC) and (QG), any solution to (3.8) must be induced by a map. Indeed, it suffices to show that (3.8) admits a unique solution, so, towards a contradiction, assume that there were distinct solutions $\pi_{1},\pi_{2}$ to (3.8). Since the objective of (3.8) is affine, it follows that $\frac{1}{2}(\pi_{1}+\pi_{2})$ is also a solution. However, if $\pi_{1},\pi_{2}$ are induced by maps, then $\frac{1}{2}(\pi_{1}+\pi_{2})$ cannot be induced by a map. This is a contradiction, hence the proof is complete. ∎
Appendix E Proofs from Section 4
@@
E.1 Proofs from Subsection 4.1
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\top}\right\|_{\textnormal{F}}\leq CD(D+\tilde{D}),$ (E.1)
@@
in (E.1) an note that $\hat{M}$ is the sample covariance matrix of $\{x_{i}\}_{i=1}^{n}$ and $M$ is the population covariance matrix of $\{y_{i}\}_{i=1}^{n}$. Letting $\bar{M}$ denote the sample covariance matrix
@@
$\|\hat{M}-M\|_{\textnormal{F}}^{2}\leq 2\|\hat{M}-\bar{M}\|_{\textnormal{F}}^{2}+2\|\bar{M}-M\|_{\textnormal{F}}^{2}.$ (E.2)
@@
where the finiteness holds by (FM). For the first term, note by (E.1) that it suffices to show $D^{2}=O_{\mathbb{P}}(\alpha_{n})$ and $\tilde{D}^{2}=O_{\mathbb{P}}(1)$; the latter follows from classical asymptotic statistics, as above, and the former holds by assumption (EBQ). This finishes the proof. ∎
@@
$\begin{split}\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{VCB}}(Z_{i})-\delta_{\mathcal{VCB}}(Z_{i})\right\|_{2}^{2}&\leq\left\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}\frac{3}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\\ &+\left\|\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}\frac{3}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\\ &+3\left\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\right\|_{2\to 2}^{2}\|\hat{\mu}\|_{2}^{2},\end{split}$ (E.3)
@@
For the third term of (E.3), it suffices to show first $\mathbb{P}(\hat{M}\preceq\hat{A})\to 1$ and then $n\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\|_{2\to 2}^{2}\|\hat{\mu}\|_{2}^{2}=O_{\mathbb{P}}(1)$ on $\{\hat{M}\preceq\hat{A}\}$; this is because $2(q-2)/q\leq 1$ for all $2<q\leq 4$, hence $O_{\mathbb{P}}(n^{-1})\leq O_{\mathbb{P}}(n^{-2(q-2)/q})$. The former follows since $\{(M,A)\in\mathbb{K}(m):M\not\prec A\}$ is an open set containing $(M,A)$ (recall (1.2)), and we have $\hat{M}\to M$ by Lemma E.1, as well as $\hat{A}\to A$ by combining (EC) with $\hat{S}\to S$ that follows from (FM) and the strong law of large numbers. To see the latter, note that $\hat{M}\preceq\hat{A}$ implies $\mathbf{t}_{\hat{M}}^{\hat{A}}\preceq I$ hence $\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\|_{2\to 2}\leq 1$. Moreover, we have $\mathbb{E}[\|Z\|^{2}]=\textnormal{tr}(A)+\textnormal{tr}(\Sigma)<\infty$ by (M) and (FM), so classical asymptotic statistics yields $\|\hat{\mu}\|_{2}^{2}=O_{\mathbb{P}}(n^{-1})$.
@@
Lastly, we consider the first term of (E.3). Indeed, we note that it suffices to show
$\left\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}=O_{\mathbb{P}}\left(\alpha_{n}\vee n^{-2(q-2)/q}\right)$ (E.4)
@@
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}=O_{\mathbb{P}}(1)$ (E.5)
@@
on the event $\{\|\hat{A}-A\|_{\textnormal{F}}\leq r\}$. Since we have $\mathbb{P}(\|\hat{M}-M\|_{\textnormal{F}}\leq r,\|\hat{A}-A\|_{\textnormal{F}}\leq r)\to 1$ by (EC) and Lemma E.1, it suffices to show that all three of
@@
are $O_{\mathbb{P}}(\alpha_{n}\vee n^{-2(q-2)/q})$ as $n\to\infty$. As we have already seen, the first is $O_{\mathbb{P}}(n^{-2(q-2)/q})$ from classical asymptotic statistics, the second is $O_{\mathbb{P}}(n^{-1})\leq O_{\mathbb{P}}(n^{-2(q-2)/q})$ by (EC), and the third is $O_{\mathbb{P}}(\alpha_{n}\vee n^{-2(q-2)/q})$ by Lemma E.1. This establishes (E.4).
To see (E.5), we simply bound:
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\leq\frac{2}{n}\sum_{i=1}^{n}\left\|\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}+\frac{2}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}.$ (E.6)
Lastly, note that the first term is $O_{\mathbb{P}}(1)$ by classical asymptotic statistics and (FM), and the second term is $O_{\mathbb{P}}(\alpha_{n})$ by another application of (EBQ). This shows that the sum is $O_{\mathbb{P}}(1)$, which establishes (E.5). This finishes the proof. ∎
E.2 Proofs from Subsection 4.2
@@
$\frac{1}{2L}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\,\textnormal{d}\bar{F}_{n}(z)$$\leq$ $\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\textnormal{d}\hat{G}_{n}(\eta)-\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\textnormal{d}\bar{G}_{n}(\eta)$ (E.7)
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}\textnormal{d}\hat{\pi}_{n}$ $=$ $\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\|\hat{\delta}_{\mathcal{B}}(z)-\eta\|^{2}\textnormal{d}\hat{\pi}_{n}(z,\eta)$ (E.8)
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\bar{\pi}_{n}$ $=$ $\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\|\delta_{\mathcal{B}}(z)-\eta\|^{2}\textnormal{d}\bar{\pi}_{n}(z,\eta)$ (E.9)
@@
Thus, equations (E.9) and (E.8) are both true. By subtracting (E.9) from (E.8), and plugging the result into (E.7) and rearranging, we conclude:
$\begin{split}\frac{1}{2L}&\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)\\ &\leq\int_{\mathbb{R}^{m}}\left(\phi^{\ast}(\eta)-\frac{1}{2}\|\eta\|^{2}\right)\textnormal{d}\hat{G}_{n}(\eta)-\int_{\mathbb{R}^{m}}\left(\phi^{\ast}(\eta)-\frac{1}{2}\|\eta\|^{2}\right)\textnormal{d}\bar{G}_{n}(\eta)\\ &\qquad-2\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)+\int\left\|\delta_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)-\int\big{\|}\hat{\delta}_{\mathcal{B}}(z)\big{\|}^{2}\textnormal{d}\bar{F}_{n}(z)\\ &\qquad+\frac{1}{2}\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}\textnormal{d}\hat{\pi}_{n}-\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\hat{\pi}_{n}\right)+\frac{1}{2}\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\hat{\pi}_{n}-\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\bar{\pi}_{n}\right).\end{split}$ (E.10)
@@
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\eta-\hat{\eta}\|^{2}\textnormal{d}\lambda_{n}(\eta,\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\nabla\phi^{\ast}(\bar{\eta})-\hat{\eta}\|^{2}\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta}\|^{2}\textnormal{d}\bar{G}_{n}(\bar{\eta})+\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\bar{\eta}-\hat{\eta}\|^{2}\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &\qquad+2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta}).\end{split}$ (E.11)
@@
because $\nabla\phi$ is an optimal transport map, and that the second term is $W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})$ by definition. Therefore, (E.11) becomes
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\eta-\hat{\eta}\|^{2}\textnormal{d}\lambda_{n}(\eta,\hat{\eta})\\ &=W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\bar{G}_{n}\right)+W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})\\ &\qquad+2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta}).\end{split}$ (E.12)

### v2

@@ Appendix F Proofs from Section 3
$\begin{cases}\textnormal{minimize}&\textnormal{Var}(\delta(Z)-\delta_{\mathcal{B}}(Z))\\ \textnormal{over}&\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}\\ \textnormal{with}&\mathbb{E}[\delta(Z)]=\mathbb{E}[\Theta]\\ \textnormal{and}&\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta).\end{cases}$ (F.1)
Now we apply a “Gaussianization trick” whereby we show that problem (F.1) has the same optimal value as the following problem:
$\begin{cases}\textnormal{minimize}&\textnormal{Var}(D-E)\\ \textnormal{over}&\textnormal{jointly Gaussian }(D,E)\\ \textnormal{with}&\textnormal{marginal distribution }D\sim\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\Theta))\\ \textnormal{and}&\textnormal{marginal distribution }E\sim\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))\\ \end{cases}$ (F.2)
To see that (F.2) $\leq$ (F.1), it suffices to find, for each $\delta$ that is feasible for (F.1), some $(D,E)$ that is feasible for (F.2) such that $\textnormal{Var}(D-E)=\textnormal{Var}(\delta(Z)-\delta_{\mathcal{B}}(Z))$. Of course, we can let
$\begin{pmatrix}D\\ E\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mathbb{E}[\Theta]\\ \mathbb{E}[\Theta]\end{pmatrix},\begin{pmatrix}\textnormal{Cov}(\Theta)&\textnormal{Cov}(\delta(Z),\delta_{\mathcal{B}}(Z))\\ \textnormal{Cov}(\delta_{\mathcal{B}}(Z),\delta(Z))&\textnormal{Cov}(\delta_{\mathcal{B}}(Z))\end{pmatrix}\right)$ (F.3)
@@
To see that (F.1) $\leq$ (F.2), it suffices to show that the optimal value of (F.2) is achieved by a coupling which corresponds to some feasible $\delta$ for (F.1) (via (F.3)). When combined with the above, this will show that such a $\delta$ is optimal for (F.1).
Note that (F.2) is indeed the (squared) 2-Wasserstein distance between the two multivariate normal distributions $\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\delta_{\mathcal{B}}(Z)))$ and $\mathcal{N}(\mathbb{E}[\Theta],\textnormal{Cov}(\Theta))$. Indeed, recall [Dowson and Landau, 1982, Olkin and Pukelsheim, 1982] that the optimal value of (F.2) is exactly given by the Bures-Wasserstein distance
@@
where $\Pi(\Sigma,\Sigma^{\prime}):=W_{2}(\mathcal{N}(0,\Sigma),\mathcal{N}(0,\Sigma^{\prime}))$, so, in particular we can use (PD) to define
@@
Proof of Theorem C.3.
First let us show that the integrand of (C.2) is well-defined for every $\pi\in\Gamma(F;G,\Psi)$. To see this, write $H$ for the second marginal of $\pi$, note that
@@
Next let us show that (C.2) admits solutions. By (C), there exists a Borel set $A\subseteq\mathbb{R}^{d}$ with $F(A)=1$ such that $c_{G}:A\times\mathbb{R}^{m}\to[0,\infty)$ is continuous. Since
@@
for all $\pi\in\Gamma(F;G)$, it follows from the Portmanteau lemma that the objective of (C.2) is weakly lower semi-continuous. Since a lower semi-continuous function on a compact set must achieve a minimizer, it only remains to show that the feasible set $\Gamma(F;G,\Psi)$ is weakly compact. To do this, note that we can write $\Gamma(F;G,\Psi)=\bigcup_{H\in\mathcal{H}}\Gamma(F;H)$ for
@@
as $r\to\infty$. This proves that $\mathcal{H}$ is tight, and hence guarantees the existence of a solution to (C.2).
Now we finish the proof, by applying the result in García-Trillos and Sen [2024, Theorem 2.4 and Remark 4.2] which shows that, under assumptions (BC) and (QG), any solution to (C.2) must be induced by a map. Indeed, it suffices to show that (C.2) admits a unique solution, so, towards a contradiction, assume that there were distinct solutions $\pi_{1},\pi_{2}$ to (C.2). Since the objective of (C.2) is affine, it follows that $\frac{1}{2}(\pi_{1}+\pi_{2})$ is also a solution. However, if $\pi_{1},\pi_{2}$ are induced by maps, then $\frac{1}{2}(\pi_{1}+\pi_{2})$ cannot be induced by a map. This is a contradiction, hence the proof is complete. ∎
Appendix G Proofs from Section 4
@@
G.1 Proofs from Subsection 4.1
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})(x_{i}-\bar{x})^{\top}-\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\top}\right\|_{\textnormal{F}}\leq CD(D+\tilde{D}),$ (G.1)
@@
in (G.1) an note that $\hat{M}$ is the sample covariance matrix of $\{x_{i}\}_{i=1}^{n}$ and $M$ is the population covariance matrix of $\{y_{i}\}_{i=1}^{n}$. Letting $\bar{M}$ denote the sample covariance matrix
@@
$\|\hat{M}-M\|_{\textnormal{F}}^{2}\leq 2\|\hat{M}-\bar{M}\|_{\textnormal{F}}^{2}+2\|\bar{M}-M\|_{\textnormal{F}}^{2}.$ (G.2)
@@
where the finiteness holds by (FM). For the first term, note by (G.1) that it suffices to show $D^{2}=O_{\mathbb{P}}(\alpha_{n})$ and $\tilde{D}^{2}=O_{\mathbb{P}}(1)$; the latter follows from classical asymptotic statistics, as above, and the former holds by assumption (EBQ). This finishes the proof. ∎
@@
$\begin{split}\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{VCB}}(Z_{i})-\delta_{\mathcal{VCB}}(Z_{i})\right\|_{2}^{2}&\leq\left\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}\frac{3}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\\ &+\left\|\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}\frac{3}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\\ &+3\left\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\right\|_{2\to 2}^{2}\|\hat{\mu}\|_{2}^{2},\end{split}$ (G.3)
@@
For the third term of (G.3), it suffices to show first $\mathbb{P}(\hat{M}\preceq\hat{A})\to 1$ and then $n\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\|_{2\to 2}^{2}\|\hat{\mu}\|_{2}^{2}=O_{\mathbb{P}}(1)$ on $\{\hat{M}\preceq\hat{A}\}$; this is because $2(q-2)/q\leq 1$ for all $2<q\leq 4$, hence $O_{\mathbb{P}}(n^{-1})\leq O_{\mathbb{P}}(n^{-2(q-2)/q})$. The former follows since $\{(M,A)\in\mathbb{K}(m):M\not\prec A\}$ is an open set containing $(M,A)$ (recall (1.2)), and we have $\hat{M}\to M$ by Lemma G.1, as well as $\hat{A}\to A$ by combining (EC) with $\hat{S}\to S$ that follows from (FM) and the strong law of large numbers. To see the latter, note that $\hat{M}\preceq\hat{A}$ implies $\mathbf{t}_{\hat{M}}^{\hat{A}}\preceq I$ hence $\|I-\mathbf{t}_{\hat{M}}^{\hat{A}}\|_{2\to 2}\leq 1$. Moreover, we have $\mathbb{E}[\|Z\|^{2}]=\textnormal{tr}(A)+\textnormal{tr}(\Sigma)<\infty$ by (M) and (FM), so classical asymptotic statistics yields $\|\hat{\mu}\|_{2}^{2}=O_{\mathbb{P}}(n^{-1})$.
@@
Lastly, we consider the first term of (G.3). Indeed, we note that it suffices to show
$\left\|\mathbf{t}_{\hat{M}}^{\hat{A}}-\mathbf{t}_{M}^{A}\right\|_{2\to 2}^{2}=O_{\mathbb{P}}\left(\alpha_{n}\vee n^{-2(q-2)/q}\right)$ (G.4)
@@
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}=O_{\mathbb{P}}(1)$ (G.5)
@@
on the event $\{\|\hat{A}-A\|_{\textnormal{F}}\leq r\}$. Since we have $\mathbb{P}(\|\hat{M}-M\|_{\textnormal{F}}\leq r,\|\hat{A}-A\|_{\textnormal{F}}\leq r)\to 1$ by (EC) and Lemma G.1, it suffices to show that all three of
@@
are $O_{\mathbb{P}}(\alpha_{n}\vee n^{-2(q-2)/q})$ as $n\to\infty$. As we have already seen, the first is $O_{\mathbb{P}}(n^{-2(q-2)/q})$ from classical asymptotic statistics, the second is $O_{\mathbb{P}}(n^{-1})\leq O_{\mathbb{P}}(n^{-2(q-2)/q})$ by (EC), and the third is $O_{\mathbb{P}}(\alpha_{n}\vee n^{-2(q-2)/q})$ by Lemma G.1. This establishes (G.4).
To see (G.5), we simply bound:
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}\leq\frac{2}{n}\sum_{i=1}^{n}\left\|\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}+\frac{2}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|_{2}^{2}.$ (G.6)
Lastly, note that the first term is $O_{\mathbb{P}}(1)$ by classical asymptotic statistics and (FM), and the second term is $O_{\mathbb{P}}(\alpha_{n})$ by another application of (EBQ). This shows that the sum is $O_{\mathbb{P}}(1)$, which establishes (G.5). This finishes the proof. ∎
G.2 Proofs from Subsection 4.2
@@
$\frac{1}{2L}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\,\textnormal{d}\bar{F}_{n}(z)$$\leq$ $\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\textnormal{d}\hat{G}_{n}(\eta)-\int_{\mathbb{R}^{m}}\phi^{\ast}(\eta)\textnormal{d}\bar{G}_{n}(\eta)$ (G.7)
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}\textnormal{d}\hat{\pi}_{n}$ $=$ $\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\|\hat{\delta}_{\mathcal{B}}(z)-\eta\|^{2}\textnormal{d}\hat{\pi}_{n}(z,\eta)$ (G.8)
@@
$\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\bar{\pi}_{n}$ $=$ $\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\|\delta_{\mathcal{B}}(z)-\eta\|^{2}\textnormal{d}\bar{\pi}_{n}(z,\eta)$ (G.9)
@@
Thus, equations (G.9) and (G.8) are both true. By subtracting (G.9) from (G.8), and plugging the result into (G.7) and rearranging, we conclude:
$\begin{split}\frac{1}{2L}&\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)\\ &\leq\int_{\mathbb{R}^{m}}\left(\phi^{\ast}(\eta)-\frac{1}{2}\|\eta\|^{2}\right)\textnormal{d}\hat{G}_{n}(\eta)-\int_{\mathbb{R}^{m}}\left(\phi^{\ast}(\eta)-\frac{1}{2}\|\eta\|^{2}\right)\textnormal{d}\bar{G}_{n}(\eta)\\ &\qquad-2\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)+\int\left\|\delta_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)-\int\big{\|}\hat{\delta}_{\mathcal{B}}(z)\big{\|}^{2}\textnormal{d}\bar{F}_{n}(z)\\ &\qquad+\frac{1}{2}\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\hat{c}_{n}\textnormal{d}\hat{\pi}_{n}-\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\hat{\pi}_{n}\right)+\frac{1}{2}\left(\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\hat{\pi}_{n}-\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}c_{G}\textnormal{d}\bar{\pi}_{n}\right).\end{split}$ (G.10)
@@
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\eta-\hat{\eta}\|^{2}\textnormal{d}\lambda_{n}(\eta,\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\nabla\phi^{\ast}(\bar{\eta})-\hat{\eta}\|^{2}\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta}\|^{2}\textnormal{d}\bar{G}_{n}(\bar{\eta})+\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\bar{\eta}-\hat{\eta}\|^{2}\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &\qquad+2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta}).\end{split}$ (G.11)
@@
because $\nabla\phi$ is an optimal transport map, and that the second term is $W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})$ by definition. Therefore, (G.11) becomes
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\eta-\hat{\eta}\|^{2}\textnormal{d}\lambda_{n}(\eta,\hat{\eta})\\ &=W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\bar{G}_{n}\right)+W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})\\ &\qquad+2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta})-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta}).\end{split}$ (G.12)